Write an equation of absolute value with x=-5 and x=-10. Some directions said to subtract one from the other and divide by 2 which is 2.5. But unsure how to figure out. Thanks
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39623)

What is the exact, word-for-word problem description.

Absolute value is Distance of a number from 0 on a number line. Absolute value is never negative. Inputs for absolute value can be positive, zero, or negative.

Answer by ikleyn(52832)

.
Although your wording is far from perfect, I can guess what you want:

|x - (-7.5)| = 2.5.

Or, which is the same,

|x + 7.5| = 2.5.



Plots y = |x + 7.5| (red) and y = 2.5 (green)